Given the following consumption function: C= 100+ 0.2y. If income is equal to R2000 consumption will be
1
Expert's answer
2019-04-11T09:28:20-0400

If

C=100+0.2y

Then y = 2000

C=100+0.2*2000=500

consumption will be R500